# SAR-Net: Multi-scale Direction-aware SAR Network via Global Information Fusion

### v2

@@ III-B1 Multi-scale Alignment Module
III-B2 Multi-level Fusion Module
Due to the typically high resolution and rich detailed information in SAR images, small objects manifest as shallow texture details, while large objects exhibit deep semantic information. Performing fusion only once cannot fully preserve the information of objects at different scales. Therefore, to enhance the model’s ability to detect multi-scale objects, we designed two Multi-level Fusion modules: the Shallow Multi-level Fusion Module (Shallow-MFM) and the Deep Multi-level Fusion Module (Deep-MFM).
@@ III-B2 Multi-level Fusion Module
Deep-MFM. Due to the advantages of transformers in handling global relationships and long-distance dependencies, and considering that large objects require more global information for accurate detection and localization, we employ a transformer fusion module in Deep-MFM. As illustrated in Fig. 4 (b), it consists of a stack of $N$ transformer blocks, where each transformer block includes Multi-Head Attention (MHA) and a Feed-Forward Network (FFN). We follow the configuration for Multi-Head Attention from [48], and this part of the process is expressed by the formula:
@@
III-B3 Multi-feature Embedding Module
Inspired by the fusion architecture with two branches in [49], to embed global information more effectively at different levels, we employ attention operations for information fusion, as illustrated in Fig. 5 Specifically, we input local information (referring to the features at the current level) and globally embedded information (generated by MFM) into two separate 1D convolutions for feature transformation. The Sigmoid function is utilized to generate attention weights from global information, which are then applied to weight the local information. Additionally, for further compensation, a residual connection is employed. This is aimed at achieving deep fusion and enhancement of global and local information. Finally, the DA Block is used for further feature extraction and fusion.
III-C Direction-aware Attention Module
Research on object detection in SAR images has shown that attention mechanisms can significantly enhance model performance. However, existing attention mechanisms often concentrate solely on positional information, frequently neglecting directional cues. This results in a lack of a method that seamlessly integrates both directional and positional information and embeds them into the attention map. In response to this, we propose the direction-aware attention module.
As shown in Fig. 6, it can be regarded as a means of enhancing the feature map’s representational capacity, obtaining direction-aware feature maps through two steps: Direction-aware Generation and Channel Attention Embedding. These steps contribute to facilitating the subsequent extraction of object features. Below, we provide a detailed description of these two steps.
III-C1 Direction-aware Generation
@@
Based on the analysis above, deformable convolution, capable of adapting to and describing object deformations well, is employed. We perform deformable convolution along one direction (either the row or column direction) on the input features, generating two directional attention maps. These attention maps are then used to modulate the input features, and this process can be expressed as:
$X_{direction}=(\sigma(DConv_{row}(X)+DConv_{column}(X)))*X,$ (7)
here, $X$ represents the input features, and $\sigma$ denotes the Sigmoid function. This bidirectional convolution attention mechanism facilitates the extraction of directional information from low-level features in two directions, thereby guiding the refinement of input features.
@@
$Z^{h}=\frac{1}{W}\sum_{i=0}^{W}X(h,i)\quad h=0,1,...,H,$ (8)
$Z^{w}=\frac{1}{H}\sum_{j=0}^{H}X(j,w)\quad w=0,1,...,W,$ (9)
@@
III-C2 Channel Attention Embedding
@@
where $\left[...\right]_{s}$ denotes the concatenation operation along the spatial dimensions, and CBR represents convolution, batch normalization, and rectified linear unit (RELU) activation function. It encodes the features with both horizontal and vertical directional information after reintegration. Subsequently, we decompose the concatenated features into $g^{h}$ and $g^{w}$ along the spatial dimension. Two $1D$ convolutions are employed to restore the feature dimensions, and finally, the weight information from the Sigmoid function is applied to the input features, resulting in:
$Y=X*\sigma\left(Conv_{h}(g^{h})\right)*\sigma(Conv_{w}(g^{w})).$ (11)
@@
III-D Loss Function
The overall loss of our network consists of two components [47]: the loss of object classification $L_{cls}$ and the loss related to bounding box regression $L_{reg}$. It can be expressed as:
@@ III-D Loss Function
we adopt the VariFocal Loss (VFL) [50] as the classification loss $L_{cls}$. VFL balances the learning signals between positive and negative samples by considering their varying levels of importance. The bounding box regression loss $L_{reg}$ comprises IOU loss and probability loss, as shown below:
@@
here, $\lambda_{iou}$ and $\lambda_{dfl}$ are hyperparameters. We choose the advanced SIOU [51] as the IOU loss $L_{iou}$. Due to the consideration of ambiguity and uncertainty in the data without introducing additional strong priors, the Distribution Focal Loss (DFL) [52] contributes to improving box localization accuracy. Therefore, we introduce an additional loss $L_{dfl}$, which is attributed to the bounding box regression loss.
IV Experiments and Analysis
To verify the superiority of SAR-Net, we perform extensive experiments on three public benchmark SAR object detection datasets, a fine-grained aircraft detection dataset SAR-AIRcraft-1.0 [53] and two ship detection datasets SSDD [54], HRSID [55]. Furthermore, we compared our proposed SAR-Net with the existing SOTA detection methods, and the performance of SAR-Net proved to be superior.
TABLE I: The Number of Instance Objects of Different Categories on the SAR-AIRcraft-1.0
@@ IV Experiments and Analysis
IV-A Dataset and Implementation Details
To most accurately evaluate the performance and generalizability of our SAR-Net framework, we conducted extensive validation experiments on a recently released fine-grained multi-class SAR aircraft object detection dataset and two commonly used SAR ship object detection datasets.
TABLE II: Ablation Study of Three Modules on Unity Compensation Mechanism
Shallow-MFM Deep-MFM MEM mAP50(%) F1-score Params(M) GFLOPs
86.5 0.813 4.64 17.80
$\checkmark$ 87.1 0.829 4.18 16.51
$\checkmark$ 85.9 0.819 4.69 15.78
$\checkmark$ 86.8 0.827 4.80 17.53
$\checkmark$ $\checkmark$ 87.8 0.835 4.87 16.82
$\checkmark$ $\checkmark$ $\checkmark$ 88.5 0.846 5.67 18.84
IV-A1 SAR-AIRcraft-1.0
@@ IV-A Dataset and Implementation Details
All images in the SAR-AIRcraft-1.0 dataset are collected from the Gaofen-3 satellite. The polarization method is single polarization, the spatial resolution is $1m$, and the imaging mode is the spotlight. Considering the size of the airport and the number of parked aircraft, the dataset mainly uses image data of three civil airports: Shanghai Hongqiao Airport, Beijing Capital Airport, and Taiwan Taoyuan Airport, including a total of $800\times 800$, $1000\times 1000$, $1200\times 1200$ and $1500\times 1500$. 4 different sizes, a total of 4,368 pictures and 16,463 aircraft object instances. The specific categories of aircraft include A220, A320/321, A330, ARJ21, Boeing737, Boeing787, other. The number of instance objects of different categories are shown in Table I, where other represents aircraft instances that do not belong to the other 6 categories. The size distribution of the aircraft object slices in the dataset spans a wide range. As shown in Fig. 7 (a), some objects are smaller than $50\times 50$, and some aircraft objects are larger than $100\times 100$. The overall object size is multi-scale. In addition, in this paper, we use the default dataset partitioning by [53] and resize all images to $800\times 800$ pixels. For the sake of fairness, we adopt the above settings when compared with other methods.
TABLE III: Research on the Effectiveness and Optimal Embedding Locations of Direction-aware Attention Module
Settings mAP50(%) F1-score Params(M) GFLOPs
Baseline 86.5 0.813 4.64 17.80
+Row Attention 87.2 0.825 4.89 18.83
+Column Attention 86.9 0.824 4.89 18.83
Behind RELU 87.4 0.831 4.82 18.69
Behind ERBlock 87.1 0.819 4.79 18.35
DAM 87.8 0.834 5.06 19.76
IV-A2 SSDD
@@
The SSDD dataset comprises data primarily sourced from Radarsat-2, TerraSAR-X, and Sentinel-1 sensors, featuring four polarization modes: HH, HV, VV, and VH. SAR image resolutions range from $1m$ to $15m$. Comprising 1,160 SAR images and 2,456 ship objects, each image contains an average of 2.12 ship objects. The height-width distribution of each ship object is illustrated in Fig. 7 (b). For standardized benchmarking, the SAR images with serial numbers ending in 1 and 9 in the SSDD dataset form the test set, ensuring an 8:2 ratio division with the training set. Consequently, the training set consists of 928 images, and the test set comprises 232 images. To facilitate fair comparisons with other methods, all images are resized to $640\times 640$ pixels in all experiments, maintaining consistent data distribution. This approach ensures a level playing field for evaluating the performance of different methods on the SSDD dataset.
TABLE IV: Collaborative Research Within the Framework, UC Represents Unity Compensation
UC Mechanism DAM mAP50(%) F1-score Params(M) GFLOPs
86.5 0.813 4.64 17.80
88.5 0.846 5.67 18.84
87.8 0.834 5.06 19.76
90.2 0.856 6.09 20.80
IV-A3 HRSID
@@
To assess the ship detection capabilities of our proposed method across different datasets, we conducted performance validation on the HRSID dataset—a large-scale, high-resolution SAR image ship object detection dataset designed for tasks such as object detection and instance segmentation. The HRSID dataset encompasses SAR images with diverse properties, including variations in sea area, polarization, sea condition, resolution, and coastal port scenarios. As depicted in Fig. 7 (c), the dataset encompasses ships at multiple scales, annotated with bounding boxes within various environments (e.g., different polarization modes, sensor types, and scenes). Panoramic images with resolutions ranging from $1m$ to $5m$ were cropped to a uniform size of $800\times 800$. In total, the HRSID dataset consists of 5,604 cropped SAR images, containing 16,951 ships. These images are distributed in a ratio of 6.5:3.5 between the training and test sets. This dataset diversity ensures comprehensive evaluation and validation of our proposed method’s performance under various conditions and scenarios.
@@
Our proposed framework is implemented in PyTorch and runs on a workstation with an NVIDIA A100 GPU. During training, the batch size was set to 32, and the initial learning rate was 0.02, reduced to 0.0002 in a cosine decay manner, and trained for a total of 300 epochs to ensure it fully converged. Standard Stochastic Gradient Descent (SGD) was used as the optimizer, and the parameters were all at default values. For input images during training, we only used mosaics augment, discarding Hue Saturation Values (HSV) and random flipping as they are not beneficial for SAR images.
TABLE V: Comparison with Different FPN Variants
NECK mAP50(%) F1-score Params(M) GFLOPs
FPN[11] 87.4 0.833 4.55 17.60
PAFPN[56] 87.2 0.828 4.82 18.13
AFPN[43] 85.2 0.807 4.67 18.19
RepBi-PAN[47] 86.5 0.813 4.64 17.80
UC Mechanism 88.5 0.846 5.67 18.84
IV-B Accuracy Metrics

### v4

@@ 3.2.1 Multi-scale alignment module
3.2.2 Multi-level fusion module
Due to the typically high resolution and rich detailed information in SAR images, small objects manifest as shallow texture details, while large objects exhibit deep semantic information. Performing fusion only once cannot fully preserve the information of objects at different scales, to augment the model’s proficiency in detecting multi-scale objects, we devised two Multi-level Fusion modules: the Shallow Multi-level Fusion Module (Shallow-MFM) and the Deep Multi-level Fusion Module (Deep-MFM).
@@ 3.2.2 Multi-level fusion module
Deep-MFM. Due to the advantages of transformers in handling global relationships and long-distance dependencies, and considering that large objects require more global information for accurate detection and localization, we employ a transformer fusion module in Deep-MFM. As depicted in Fig. 4 (b), it comprises a stack of $N$ transformer blocks, with each block containing Multi-Head Attention (MHA) and a Feed-Forward Network (FFN). We follow the configuration for Multi-Head Attention from [24], and this part of the process is expressed by the formula:
@@
3.2.3 Multi-feature embedding module
Inspired by the fusion architecture with two branches, to embed global information more effectively at different levels, we employ attention operations for information fusion, as illustrated in Fig. 5. In particular, we feed local information (pertaining to features at the operating level) and globally embedded information (generated by MFM) into two separate 1D convolutions for feature transformation. The Sigmoid function is utilized to generate attention weights from global information, which are then applied to weight the local information. Additionally, for further compensation, a residual connection is employed. This is aimed at achieving deep fusion and enhancement of global and local information. Finally, the DA Block is used for further feature extraction and fusion.
3.3 Direction-aware attention module
Research on object detection in SAR images has demonstrated that attention mechanisms can notably enhance model performance. However, existing attention mechanisms often concentrate solely on positional information, frequently neglecting directional cues. This results in a lack of a method that seamlessly integrates both directional and positional information and embeds them into the attention map. In response to this, we introduce the direction-aware attention module.
As depicted in Fig. 6, it can be regarded as a means of enhancing the feature map’s representational capacity, obtaining direction-aware feature maps through two steps: Direction-aware Generation and Channel Attention Embedding. These steps contribute to facilitating the subsequent extraction of object features. Below, we provide a detailed description of these two steps.
@@ 3.3 Direction-aware attention module
3.3.1 Direction-aware generation
@@
Based on the analysis above, deformable convolution, capable of adapting to and describing object deformations well, is employed. We perform deformable convolution along one direction (either the row or column direction) on the input features, generating two directional attention maps. These attention maps are then used to modulate the input features, and it is expressed as:
$X_{direction}=(\varphi(DConv_{row}(X)+DConv_{column}(X)))\otimes X,$ (7)
here, $X$ represents the input features, $\varphi$ represents the Sigmoid function, $\otimes$ represents element-wise matrix multiplication, as the dimensions of the matrices before and after the operation are equal. This bidirectional convolution attention mechanism facilitates the extraction of directional information from low-level features in two directions, thereby guiding the refinement of input features.
@@
$Z^{h}=\frac{1}{W}\sum_{i=0}^{W}X_{direction}(h,i)\quad h=0,1,...,H,$ (8)
$Z^{w}=\frac{1}{H}\sum_{j=0}^{H}X_{direction}(j,w)\quad w=0,1,...,W,$ (9)
@@
3.3.2 Channel attention embedding
@@
where $\left[...\right]_{s}$ denotes the concatenation operation along the spatial dimensions, and CBR represents the sequence of convolution, batch normalization, and RELU activation function. It encodes the features with both horizontal and vertical directional information after reintegration. Subsequently, we decompose the concatenated features into $g^{h}$ and $g^{w}$ along the spatial dimension. Two $1D$ convolutions are employed to restore the feature dimensions, and finally, the weight information from the Sigmoid function is applied to the input features, resulting in:
$Y=X*\varphi\left(Conv_{h}(g^{h})\right)*\varphi(Conv_{w}(g^{w})),$ (11)
where * represents tensor product.
@@
3.4 Loss function
The overall loss of our network consists of two components: the loss of object classification $L_{cls}$ and the loss related to bounding box regression $L_{reg}$. It can be expressed as:
@@ 3.4 Loss function
we adopt the VariFocal Loss (VFL) as the classification loss $L_{cls}$. VFL balances the learning signals between positive and negative samples by considering their varying levels of importance. The bounding box regression loss $L_{reg}$ comprises IOU loss and probability loss, as shown below:
@@
here, $\lambda_{iou}$ and $\lambda_{dfl}$ are hyperparameters. We choose the advanced SIOU as the IOU loss $L_{iou}$. Due to the consideration of ambiguity and uncertainty in the data without introducing additional strong priors, the Distribution Focal Loss (DFL) contributes to improving box localization accuracy. Additionally, we add a necessary loss $L_{dfl}$, which accounts for the bounding box regression loss.
4 Experiments and analysis
To verify the superiority of SAR-Net, we perform extensive experiments on three public benchmark SAR object detection datasets, a fine-grained aircraft detection dataset SAR-AIRcraft-1.0 [25] and two ship detection datasets SSDD [26], HRSID [27]. Furthermore, we benchmarked our proposed SAR-Net against SOTA detection methods, and the results demonstrated the superior performance of SAR-Net.
4.1 Dataset and implementation details
To meticulously assess the performance and generalizability of SAR-Net framework, we executed a lot of verification experiments on a recently released fine-grained multi-class SAR aircraft object detection dataset and two commonly used SAR ship object detection datasets.
Table 1: The total of instance objects of different categories on the SAR-AIRcraft-1.0.
@@ 4.1 Dataset and implementation details
4.1.1 SAR-AIRcraft-1.0
@@
Images acquired from the Gaofen-3 satellite are screened to form the SAR-AIRcraft-1.0 dataset. The polarization method is single polarization, with a spatial resolution of $1m$, and utilizing the spotlight imaging mode. Considering the size of the airport and the number of parked aircraft, the dataset mainly uses image data of three civil airports: Shanghai Hongqiao Airport, Beijing Capital Airport, and Taiwan Taoyuan Airport, including a total of $800\times 800$, $1000\times 1000$, $1200\times 1200$ and $1500\times 1500$. 4 different sizes, a total of 4,368 pictures and 16,463 aircraft object instances. The specific categories of aircraft include A220, A320/321, A330, ARJ21, Boeing737, Boeing787, other. The total of instance objects of different categories is shown in Table 1, where other represents aircraft instances that do not belong to the other 6 categories. The size distribution of the aircraft object slices in the dataset spans a wide range. As shown in Fig. 7 (a), some objects are smaller than $50\times 50$, and some aircraft objects are larger than $100\times 100$. The overall object size is multi-scale. In addition, in this paper, we use the default dataset partitioning by [25] and resize all images to $800\times 800$ pixels. To ensure fairness, we adopt the above settings when compared with other methods.
4.1.2 SSDD
@@
The SSDD dataset comprises data primarily sourced from three sensors. Its image resolution exhibits a wide range of variations. Comprising 1,160 SAR images and 2,456 ship objects, on average, each image contains 2.12 ship objects. The sizes of each ship object are shown in Fig. 7 (b). For standardized benchmarking, the images with numbers ending in 1 and 9 in the dataset form the testset, ensuring an 8:2 ratio division with the training set. To facilitate fair comparisons with other methods, all images are resized to $640\times 640$ pixels in all experiments, maintaining consistent data distribution.
4.1.3 HRSID
@@
To further assess the ship detection capabilities of SAR-Net, we conducted experiments on HRSID, which is a large-scale, high-resolution SAR images dataset designed for ship object detection tasks. The HRSID dataset encompasses SAR images with diverse properties, including variations in sea area, polarization, sea condition, resolution, and coastal port scenarios. As depicted in Fig. 7 (c), the dataset encompasses objects at a wide range of scales, annotated with bounding boxes within various environments. Panoramic images with resolutions ranging from $1m$ to $5m$ are cropped to a uniform size of $800\times 800$. In total, the HRSID dataset consists of 5,604 cropped SAR images, containing 16,951 ships. These images are distributed in an 6.5:3.5 ratio division with the training set and test set. This dataset diversity ensures comprehensive evaluation and validation of our proposed method’s performance under various conditions and scenarios.
@@
Our proposed framework has been conducted using PyTorch and is executed on a workstation equipped with an NVIDIA A100 GPU. During the training phase, we set the batch size to 32, and the learning rate is 0.02, which gradually reduces to 0.0002 following a cosine decay schedule, and train for a total of 300 epochs to ensure it fully converged. Standard Stochastic Gradient Descent (SGD) is used as the optimizer, and the parameters are all at default values. For input images during training, we only used mosaics augment, discarding Hue Saturation Values (HSV) and random flipping as they are not beneficial for SAR images.
4.2 Accuracy metrics
@@
4.3 Ablation study
To comprehensively and detailedly analyze the effectiveness of SAR-Net, we executed extensive ablation experiments with different framework designs and internal modules, and all experiments were performed on the SAR-AIRcraft-1.0 dataset. The best-performing results are highlighted in bold.
Table 2: Ablation study of three modules on unity compensation mechanism with MAM using concatenation operation.
MAM S-MFM D-MFM MEM mAP50(%) F1-score(%) Params(M) GFLOPs
86.5 81.3 4.6 17.8
$\checkmark$ 85.9 81.9 4.0 15.5
$\checkmark$ $\checkmark$ 87.1 82.9 4.2 16.5
$\checkmark$ $\checkmark$ 87.0 82.5 4.7 15.8
$\checkmark$ $\checkmark$ 86.8 82.7 4.8 17.5
$\checkmark$ $\checkmark$ $\checkmark$ 87.5 83.3 5.0 18.5
$\checkmark$ $\checkmark$ $\checkmark$ 87.3 82.9 5.4 17.8
$\checkmark$ $\checkmark$ $\checkmark$ 87.8 83.5 4.9 16.8
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 88.5 84.6 5.6 18.8
Table 3: Ablation study of three modules on unity compensation mechanism with MAM using addition operation.
MAM S-MFM D-MFM MEM mAP50(%) F1-score(%) Params(M) GFLOPs
86.5 81.3 4.6 17.8
$\checkmark$ 84.2 80.7 4.1 15.8
$\checkmark$ $\checkmark$ 86.8 82.3 4.5 16.8
$\checkmark$ $\checkmark$ 86.8 82.2 5.0 16.1
$\checkmark$ $\checkmark$ 86.6 82.3 5.1 17.7
$\checkmark$ $\checkmark$ $\checkmark$ 87.2 83.0 5.3 18.7
$\checkmark$ $\checkmark$ $\checkmark$ 86.9 82.6 5.6 18.1
$\checkmark$ $\checkmark$ $\checkmark$ 87.1 82.5 5.2 17.0
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 88.0 84.4 5.9 19.1
4.3.1 Impact of unity compensation mechanism
To study the effectiveness of each module in the unity compensation mechanism and verify the correctness of uniting deep and shallow layers’ global information, we executed a lot of ablation experiments on four of the modules and measured the performance with mAP50 and F1-score, a common indicator in object detection. In addition, Parameter sizes and GFLOPs are also shown in Table 2 and Table 3. We use YOLOv6n [23] as the baseline for ablation experiments to embed the modules of our unity compensation mechanism one by one. The feature alignment process is included in all experiments except the baseline. In Table 2, we conducted MAM using concatenation operation, where S-MFM represents Shallow-MFM and D represents Deep-MFM. It can be seen that only when the MAM module is used, the accuracy is lower than the baseline. This is because it integrates semantic information and detailed texture information from different layers, but it lacks further fusion of object-specific information. In addition, other experiments have achieved performance improvements without a substantial increase in computational overhead, which proves the efficacy of our global information fusion idea. In particular, when the MEM module is not used, we use simple addition to compensate for the original features.
Figure 8: Layercam visualization of different modules on Unity Compensation Mechanism.
To explore a new alignment pattern, we conducted MAM using addition operation. It is worth noting that both concatenation and addition operations require aligning features from each layer to have the same dimensions. As can be seen from Table 3, the overall trends of the other modules are similar to those in Table 2. However, due to the addition operation combining rich texture and semantic information directly from different layers, it made the information redundant and resulted in difficulties in subsequent fusion compensation operations, leading to information loss and performance degradation. Additionally, features from different layers have varying channel numbers, requiring channel alignment during the addition operation, which increases parameter and computation complexity.
To visually demonstrate the progressive effect of modules within the Unity Compensation Mechanism, we utilized Layercam for visualizing feature maps. As depicted in Fig. 8, the MAM aligns features from different layers but does not perform fusion operations, resulting in a chaotic feature map with scattered focus points across the entire image. However, it still retains features from different hierarchical levels for subsequent fusion compensation operations. After passing through Shallow-MFM, detailed texture information is fused and compensated back to the original features, enabling the model to focus on each object with texture information. Subsequently, with Deep-MFM, semantic information is fused and compensated onto the texture information, allowing the model to focus on object areas without losing texture information. Finally, the MEM embeds all features, further narrowing down the focus areas, which aids in model detection. The intuitive visualization results validate that the Unity Compensation Mechanism enhances feature integration while maximizing the retention of semantic and texture information.
Table 4: Research on the effectiveness and optimal embedding locations of direction-aware attention module.
Settings mAP50(%) F1-score(%) Params(M) GFLOPs
Baseline 86.5 81.3 4.6 17.8
+Row Attention 87.2 82.5 4.9 18.8
+Column Attention 86.9 82.4 4.9 18.8
DAM w/o CAE 87.4 83.3 4.9 19.1
DAM w/o DG 86.8 82.4 4.7 18.4
Behind RELU 87.4 83.1 4.8 18.6
Behind ERBlock 87.1 81.9 4.8 18.3
DAM 87.8 83.4 5.0 19.7
4.3.2 Impact of direction-aware attention module
To prove the effectiveness of the proposed direction-aware attention module, we executed a series of detailed experiments, as shown in Table 4. We split the attention modules in two directions and embedded them in the YOLOv6n backbone separately, both received performance improvements. To explore the roles of the two modules within DAM, we conducted experiments by using only one module at a time on the baseline. Here, w/o CAE represents without Channel Attention Embedding, and w/o DG represents without Direction-aware Generation. Both experiments showed improvement compared to the baseline. However, due to the absence of crucial bidirectional feature vector extraction, DAM w/o DG showed minimal performance enhancement, merely reallocating attention among channels. Equally noteworthy, after generating the direction-aware attention map, omitting the embedding of the attention map among channels also resulted in performance degradation. Therefore, the collaboration of these two modules enables the model to concentrate more effectively on the directional areas of objects.
In addition, we tried to embed DAM in different positions of the backbone. Except for the embedding position shown in Fig. 3, we also tried to move the DAM behind the RELU and each ERBlock, and obtained 0.9% and 0.6% performance improvement respectively with a small increase in parameters. Exhaustive experiments allowed us to determine the most suitable embedding position. These experimental results show that the acquisition of direction-aware information is significantly helpful for SAR images containing directional objects.
Table 5: Collaborative research within the framework, UC represents unity compensation.
UC Mechanism DAM mAP50(%) F1-score(%) Params(M) GFLOPs
86.5 81.3 4.6 17.8
88.5 84.6 5.6 18.8
87.8 83.4 5.0 19.7
90.2 85.6 6.0 20.8
4.3.3 Collaboration of the framework
The effectiveness of the components that make up SAR-Net has been demonstrated, and to further explore the synergy between them, we combine them and find the desired results. For the unity compensation mechanism, the backbone that can extract direction-aware information gives full play to the advantages of global fusion and enriches the information of multi-scale features. For the DAM-embedded backbone, the unity compensation mechanism retains detailed information and enhances the efficiency of information fusion and transmission. As shown in Table 5, after combining the two, their collaborative complementarity is fully utilized and performance is further improved by 0.4%.
Table 6: Comparison with different FPN variants.
NECK mAP50(%) F1-score(%) Params(M) GFLOPs
FPN[3] 87.4 83.3 4.5 17.6
PAFPN[28] 87.2 82.8 4.8 18.1
AFPN[21] 85.2 80.7 4.6 18.1
RepBi-PAN[23] 86.5 81.3 4.6 17.8
UC Mechanism 88.5 84.6 5.6 18.8
4.3.4 Comparison of different FPN architecture
We choose the classic FPN structure and its latest variants AFPN [21] and RepBi-PAN [23] for comparison. We conduct the experiments based on YOLOv6n and the results are illustrated in Table 6. A significant accuracy disparity exists despite the parameters being similar. As we discuss in the section 1, sequential local fusion will lead to information loss, especially for SAR images with large object scale variants. Top-down and bottom-up structures will lose a lot of detailed information and semantic information before detection. Global information fusion and compensation preserve the object’s multi-scale information to the maximum extent, thereby enabling the unity compensation mechanism to deliver optimal performance.